Given the measured cell potential Ecel, is -0.3657 Vat 25 °C in the following cell, calculate the H+ concentration Pt (s) H2 g, 0.801 atm)IH (aq, M)IICd2 (aq, 1.00 M)lcd (s) The balanced reduction half reactions for the cell, and their respective standard reduction potential values, E are as follows 2H (aq)+ 2e E 0.00 V E 0.403 V Number M.
Answer #1

first reactlon, H2(g) ----->2H+ +2e- , Eo=0 V (1)

Cd+2+2e- -------->Cd, Eo=-0.403V (2)

Adding Eq.1 and Eq.2 gives

H2+Cd+2 -------->2H+ + Cd, Eo= -0.403V

E= Eo-(0.0591/2)*logQ

-0.3657= -0.403 -(0.0591/2)* logQ, Q= [H+]2/ [H2][Cd+2]

0.0373= -0.02955* logQ

logQ= -1.26

Q= 0.055 = [H+]2/ (0.801*1)

[H+]2= 0.055*0.801 , [H+] =0.21 M